The domain of the function f(x) = 2x is all Real numbers but is the Range all Real numbers? How would you get the number 3 in the Range?
Answer by Jk22(389)

f(x)=3=2x => x=3/2 : three halves gives as output 3
3/2*2=3/2+3/2=(3+3)/2=6/2=3